OpenStudy (anonymous):

I'm in Algebra 2 and we have to graph absolute value functions. Can someone show me the steps to making the table to solve them. An example is y=(the absolute value of) 3x +1(absolute value ends) +1?

jimthompson5910 (jim_thompson5910):

so you have this right? y = |3x+1| + 1

OpenStudy (anonymous):

Yes and I remember my teacher saying that in the middle of the table we have to make is the number that makes the absolute value zero, but I don't know what makes it zero

jimthompson5910 (jim_thompson5910):

let's plug in various values of x to see what we get

jimthompson5910 (jim_thompson5910):

so when x = 0, we will have y = |3x+1| + 1 y = |3(0)+1| + 1 y = |0+1| + 1 y = |1| + 1 y = 1+1 y = 2 That means we have the row of "0 2" in the table 

jimthompson5910 (jim_thompson5910):

let's repeat that with x = 1 y = |3x+1| + 1 y = |3(1)+1| + 1 y = |3+1| + 1 y = |4| + 1 y = 4+1 y = 5 That means we have the row of "1 5" in the table  hopefully you're seeing how all this works?